# Quadratic Gaussian Splatting for Efficient and Detailed Surface Reconstruction

### v1

@@ Abstract
Recently, 3D Gaussian Splatting (3DGS) has attracted attention for its superior rendering quality and speed over Neural Radiance Fields (NeRF). To address 3DGS’s limitations in surface representation, 2D Gaussian Splatting (2DGS) introduced disks as scene primitives to model and reconstruct geometries from multi-view images, offering view-consistent geometry. However, the disk’s first-order linear approximation often leads to over-smoothed results. We propose Quadratic Gaussian Splatting (QGS), a novel method that replaces disks with quadric surfaces, enhancing geometric fitting. QGS defines Gaussian distributions in non-Euclidean space, allowing primitives to capture more complex textures. As a second-order surface approximation, QGS also renders spatial curvature to guide the normal consistency term, to effectively reduce over-smoothing. Moreover, QGS is a generalized version of 2DGS that achieves more accurate and detailed reconstructions, as verified by experiments on DTU and TNT, demonstrating its effectiveness in surpassing most state-of-the-art methods in geometry reconstruction. Our code will be released as open source. The project is available at https://quadraticgs.github.io/QGS
Figure 1: Quadratic Gaussian Splatting, a novel quadric-based approach for recovering high-precision real-world geometry from multi-view RGB images. In the Courthouse scene, our method achieves multi-view consistent normals, depth, and curvature, producing high-quality meshes and realistic images. Compared to 2DGS, QGS more accurately models the building’s contours.
@@ 1 Introduction
Surface reconstruction and realistic novel view synthesis (NVS) are critical tasks in computer graphics and vision. Their goal is to recover dense geometric structures from images captured from different viewpoints and render photorealistic images. Recently, 3D Gaussian Splatting (3DGS) has surpassed Neural Radiance Fileds (NeRF) based methods [2, 23, 24] in both rendering quality and speed by combining traditional splatting techniques with end-to-end optimization. Subsequently, GS-based methods have seen rapid advancements in dynamic reconstruction [13, 8], editing [6, 29], and large-scale scene reconstruction [20, 21].
However, vanilla splatting ignores the Gaussian contribution along the z-axis and uses approximation [39, 16]. As a result, it can render high-quality, multi-view consistent textures even at offsets from the scene surface, making it challenging to capture accurate geometry. In subsequent work [35, 19, 12, 4, 36], ray-splat intersections are computed in 3D to retrieve Gaussian weights, a key factor for accurate scene geometry recovery. In methods with Gaussian ellipsoids as primitives [35, 36], ray-splat intersections depend on the view direction, limiting normal consistency across views. In contrast, methods using Gaussian disks as primitives [12, 19] can ensure multi-view consistent geometry due to their surface-based nature, facilitating the incorporation of multi-view geometric consistency [4, 19]. However, disks, as a first-order fit of geometry, often result in overly smooth reconstructions, as shown in the Fig 2.
Earlier works [26, 40, 31] demonstrated that surfel elements can effectively represent complex scene geometry, such as atomic models. Building on this, 2DGS [12] combined Gaussian distributions with disk elements to achieve state-of-the-art (SOTA) reconstruction, inspiring our approach. In this paper, we propose a novel surface-based rendering method, which we call Quadratic Gaussian Splatting (QGS), for 3D scene reconstruction and novel view synthesis. This method enhances the geometric fitting of scene primitives through higher-order representation, enabling the extraction of accurate surface mesh from the quadric model.
Unlike previous splatting techniques operating in Euclidean space, such as 3DGS or 2DGS, our QGS defines the Gaussian distribution on a quadratic paraboloid, which can continuously transition between convex and concave forms, offering greater geometric fitting flexibility. The key technique in using paraboloids for splatting lies in establishing Gaussian distributions based on geodesic distances in non-Euclidean space, which allows the distribution’s energy to be concentrated on the surface, to effectively capture complex geometric textures. Compared to the first-order approximation in 2DGS, QGS provides a second-order fit to the scene geometry. As a result, QGS not only produces multi-view consistent normals and depth but also provides curvature information, which represents surface bending and helps guide single-view normal consistency supervision. Furthermore, we observe that the volumetric rendering order used by most GS methods causes popping artifacts, which impact both novel view synthesis and geometric reconstruction. To address this, we adopted the sorting criteria from StopThePop [25], restructuring it to accommodate quadric surface structures, thereby improving the centroid sorting used in 2DGS to better handle complex intersections of quadratic surfaces. Finally, QGS is able to extract high-quality and detail-rich mesh models, as shown in the Fig 1.
In summary, the main contributions of our work are as follows:
  • We propose QGS, a novel and efficient differentiable representation that uses quadric surfaces as scene primitives, offering stronger geometric fitting capabilities.
  • We propose the first Gaussian Splatting work to introduce geodesic distance to establish Gaussian distributions on surfaces, enabling primitives to fit more complex textures.
  • With stricter depth sorting in the higher-degree-of-freedom QGS, our method achieves SOTA geometric reconstruction and enhanced rendering quality.
2 Related work
2.1 Novel View Synthesis
Novel view synthesis has seen rapid development in recent years, especially since the introduction of NeRF [23]. NeRF encodes scene geometry and view-dependent appearance using a multi-layer perceptron (MLP) and optimizes it end-to-end through volumetric rendering to produce realistic images. Subsequent methods introduced various improvements. Mip-NeRF [1], Mip-NeRF 360 [2], and Zip-NeRF [3] addressed NeRF’s aliasing issues by introducing new sampling strategies. I-NGP [24] and DVGO [28] significantly accelerated NeRF’s training and rendering speeds by encoding scenes into feature grids instead of MLPs.
Recently, 3DGS [16] has gained attention for its more realistic rendering capabilities and real-time rendering speed, surpassing NeRF in many cases [34, 25, 22, 38]. Mip-Splatting [34] introduced a 3D smoothing filter to eliminate high-frequency artifacts during rendering. StopThePop [25] proposed per-tile and per-pixel resorting to resolve multi-view geometric inconsistencies. Scaffold-GS [22] effectively reduced redundant Gaussians and improved rendering quality by constructing anchor points to distribute local 3D Gaussians.
2.2 Neural Surface Reconstruction
Compared to traditional multi-view geometric reconstruction, neural volumetric rendering methods often produce smoother and more complete results. NeuS [30] and VolSDF [32] introduced signed distance fields (SDF) to describe scene geometry based on NeRF [23]. Geo-NeuS [9] and NeuralWarp [10] improved reconstruction by incorporating multi-view consistency. Neuralangelo [18] reduced MLP reliance by using feature grids and numerical gradients to improve geometric reconstruction quality.
However, neural rendering-based surface reconstruction generally requires several hours to converge using only images. Additionally, due to the structured implicit fields, accessing and editing scene geometry is not straightforward.
2.3 Gaussian Splatting Surface Reconstruction
With the rapid development of 3DGS across various fields, Gaussian splatting methods have also made significant progress in surface reconstruction. SuGaR [11] introduced regularization terms to encourage Gaussians to fit the scene surface by building volumetric density fields from Gaussian ellipsoids and using Poisson reconstruction [15] to obtain a mesh. GSDF [33] and NeuSG [5] combined 3D Gaussian ellipsoids with SDF, achieving high-quality geometric reconstructions and rendering results. Rade-GS [36], GOF [35], and PGSR [4] computed ray-Gaussian ellipsoid intersections to obtain unbiased depth and used the normal consistency supervision from 2DGS [12] to achieve state-of-the-art reconstruction results. 2DGS flattened Gaussian ellipsoids into Gaussian disks to better align Gaussian primitives with surfaces and introduced two regularization losses to provide additional geometric constraints. MVG-splatting [19] extended 2DGS by introducing multi-view consistency constraints, enabling more complex surface reconstructions. In this paper, we propose Quadratic Gaussian Splatting (QGS), a new surface representation based on quadric surfaces designed to improve the local geometric fitting capability of primitives. We further establish Gaussian distributions on quadric, enabling end-to-end optimization. Lastly, we address the volume rendering order issue in 2DGS by incorporating the sorting method from StopThePop [25], achieving precise geometric reconstructions and high-quality rendering results.
3 Method
3.1 Preliminary
Kerbl et al. [16] proposed representing a scene using 3D Gaussian ellipsoids as primitives and render images using differentiable volume splatting. The shape and orientation of the ellipsoids are controlled by the scale $\mathbf{S}$ and pose $\mathbf{R}$ at their location $\mathbf{p}_{k}$. Together, $\mathbf{R}$, $\mathbf{S}$, and $\mathbf{p}_{k}$ define a metric space, within which a Gaussian distribution is formulated:
$G(\mathbf{x})=\exp(-\frac{1}{2}(\mathbf{x}-\mathbf{p}_{k})^{T}\mathbf{\Sigma}^{-1}(\mathbf{x}-\mathbf{p}_{k}))$ (1)
Where the covariance matrix $\mathbf{\Sigma}=\mathbf{RSS^{T}R^{T}}$. To render the 3D Gaussian onto an image, each Gaussian distribution is first transformed into the camera space using the world-to-camera transformation $\mathbf{W}$. It is then mapped into ray space, where the viewing direction is aligned with the coordinate axis, through a local affine transformation $\mathbf{J}$ [39].
$\mathbf{\Sigma}^{\prime}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W}^{T}\mathbf{J}^{T}$ (2)
Under orthogonal projection, we skip the third row and column of $\mathbf{\Sigma}^{\prime}$, yielding a 2×2 covariance matrix $\mathbf{\Sigma}^{2D}$ to represent the 2D Gaussian $G^{2D}$. Subsequently, 3DGS employs volume rendering to integrate the 2D Gaussian.
Here, $\alpha_{i}$ represents the opacity, $c_{i}$ denotes the color of each Gaussian primitive, and $\mathbf{p}$ is the pixel coordinate. Finally, each Gaussian primitive is optimized by minimizing the photometric loss.
GS mesh reconstruction. Vanilla 3DGS [16] can render high-quality images, but it yields suboptimal results for scene geometry reconstruction due to a lack of multi-view consistency in its splatting process. In subsequent geometric reconstruction methods, volumetric approaches like GOF [35] and RadeGS [36] employ ray-splat intersection techniques, achieving state-of-the-art reconstruction quality but limiting the consistency of normal and depth. In contrast, 2DGS [12] defines the 2D Gaussian distribution within a planar disk, which inherently provides consistent normals and depths across multiple views. But the disk is only a first-order approximation of the surface, which often leads to overly smooth reconstruction results in 2DGS, as shown in Fig 2.
@@ 3.1 Preliminary
Figure 2: Demo and mesh reconstruction comparison on Mip-NeRF 360 dataset. 3D Gaussian ellipsoids struggle to fit precisely onto surfaces, resulting in coarse and incomplete meshes. Planes cannot fully capture high-curvature regions, leading to over-smoothed reconstructions. In contrast, quadric surfaces effectively capture geometric edge regions, achieving detailed reconstruction results.
@@ 3.2 Quadratic Gaussian splatting
3.2.1 Quadratic Gaussian Model
Quadratic Model. Given a homogeneous coordinate $\mathbf{x}=[x,y,z,1]^{T}\in\mathbb{R}^{4}$, a quadric surface can be defined as the solution set to the following equation:
@@ 3.2.1 Quadratic Gaussian Model
Similar to [26], we apply congruent diagonalization to transform the above equation into its canonical form:
$\mathbf{Q}=\mathbf{T}^{-T}\mathbf{D}\mathbf{T}^{-1},\mathrm{with}$ $\mathbf{D}\;\mathrm{diagonal},d_{ii}\in\{0,\pm 1\}$ (5)
$\mathbf{T}=\left[\begin{matrix}\mathbf{u}&\mathbf{v}&\mathbf{w}&\mathbf{c}\\ 0&0&0&1\end{matrix}\right]$
Here, $\mathbf{c}$ denotes the position of the quadric, and $[\mathbf{u},\mathbf{v},\mathbf{w}]$ can be decomposed into $\mathbf{RS}$, where $\mathbf{R}=[\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}]$ and $\mathbf{S}=\mathrm{diag}(s_{1},s_{2},s_{3})$, which denotes the orientation and scale of the quadric in the object space. The matrix $\mathbf{D}$ defines the surface shape: $\mathbf{D}=\mathrm{diag}(1,1,1,1)$ yields an ellipsoid, while $\mathbf{D}=\mathrm{diag}(1,0,0,0)$ produces a plane.
To compute the Gaussian weight at any surface point, we first define a measure on the surface to establish the Gaussian distribution. To concentrate the Gaussian energy on the surface, we use geodesic length [27] as the metric: the geodesic distance between two surface points is the shortest path along the surface, as illustrated by the red line in Fig 3.
However, not all geodesics have closed-form solutions. When $\mathbf{Q}$ is an ellipsoid or hyperboloid, computing the geodesic length typically requires numerical methods. Therefore, we focus only on the case of a paraboloid:
$f(x,y,z)$ $=\mathbf{x}^{T}\left[\begin{matrix}\mathbf{R}&\mathbf{c}\\ 0&1\end{matrix}\right]^{-T}\bar{D}\left[\begin{matrix}\mathbf{R}&\mathbf{c}\\ 0&1\end{matrix}\right]^{-1}\mathbf{x}$ (6)
@@
Gaussian Distribution on Quadric. We will now describe how to define a Gaussian distribution on a paraboloid. First, we need to define a measure on the quadric. Paraboloid (Equation 7) could be expressed in explicit form:
@@
$\vspace{-5pt}\begin{split}\hat{z}(\theta,\rho)=\;&s_{3}(\frac{\mathrm{sign}(s_{1})\cos^{2}\theta}{s_{1}^{2}}+\frac{\mathrm{sign}(s_{2})\sin^{2}\theta}{s_{2}^{2}})\rho^{2}\\ =\;&a(\theta)\rho^{2}\end{split}\vspace{-5pt}$ (9)

### v3

# Quadratic Gaussian Splatting: High Quality Surface Reconstruction with Second-order Geometric Primitives

@@ Abstract
We propose Quadratic Gaussian Splatting (QGS), a novel representation that replaces static primitives with deformable quadric surfaces (e.g., ellipse, paraboloids) to capture intricate geometry. Unlike prior works that rely on Euclidean distance for primitive density modeling—a metric misaligned with surface geometry under deformation—QGS introduces geodesic distance-based density distributions. This innovation ensures that density weights adapt intrinsically to the primitive curvature, preserving consistency during shape changes (e.g., from planar disks to curved paraboloids). By solving geodesic distances in closed form on quadric surfaces, QGS enables surface-aware splatting, where a single primitive can represent complex curvature that previously required dozens of planar surfels, potentially reducing memory usage while maintaining efficient rendering via fast ray-quadric intersection. Experiments on DTU, Tanks and Temples, and MipNeRF360 datasets demonstrate state-of-the-art surface reconstruction, with QGS reducing geometric error (chamfer distance) by 33% over 2DGS and 27% over GOF on the DTU dataset. Crucially, QGS retains competitive appearance quality, bridging the gap between geometric precision and visual fidelity for applications like robotics and immersive reality.
GT
2DGS
PGSR
QGS (8K)
Figure 1: Our QGS model employs quadric surfaces to recover high-precision geometry from multi-view RGB images, reducing point count while enhancing accuracy. To isolate geometry from texture, we evaluate QGS on a weakly textured, geometry-rich subset of the Google Scanned Object dataset [9]. The results suggest that QGS, under different gradient threshold settings (QGS-x), consistently outperforms 2DGS [16] and achieves comparable quality to PGSR [5] with fewer points.
@@ 1 Introduction
The pursuit of photorealistic view synthesis and geometry reconstruction is a central focus in academia and industry, with applications spanning virtual reality, film-making, and autonomous driving. Recent point-based methods, such as 3D Gaussians [21, 50], have proven highly effective, delivering view synthesis quality comparable to Neural Radiance Fields (NeRF)-based approaches while preserving superior geometry [16, 46] and offering significantly faster performance. Since then, these methods have been extended to dynamic reconstruction [17, 10, 27], scene editing [7, 41, 40, 54], and large-scale scene reconstruction [26, 28, 44, 29].
However, due to their approximation-based nature, 3D Gaussian methods often face challenges like high memory consumption, blurriness, and geometric inaccuracies. To address these limitations, recent efforts have focused on “primitive-level modifications.” Methods like GES [14] generalize density distributions to reduce point counts, while others employ sharp kernels (e.g., clipped Gaussian [23], radial [18], or convex kernels [15]) to capture high-frequency appearance details. Assigning texture maps to planar primitives has also enabled more efficient scene representations [34, 39, 36, 4, 43]. Despite these advances, such modifications mainly target high-quality appearance reconstruction, leaving detailed geometry reconstruction largely unaddressed. Several approaches adopt alternative kernel shapes, such as flat ellipsoids or planar disks (e.g., surfels) [46, 16, 8] to improve surface alignment. Yet, as a first-order geometric approximation, planar disks lack sufficient expressiveness; For example, highly curved surfaces requires dense primitive sampling, creating memory bottlenecks and rendering latency.
In this paper, we extend the representations of ellipsoid [21] and elliptical disks to a more general quadratic framework. Specifically, we define primitive boundaries using an implicit function $f(x,y,z)=0$ where $f(x,y,z)<0$ for points inside the primitive and $f(x,y,z)>0$ otherwise. By adopting a quadratic form $f(x,y,z)=(x,y,z,1)^{\top}\mathbf{Q}(x,y,z,1)=0$, where $\mathbf{Q}\in\mathbb{R}^{4\times 4}$ is a conic matrix, we allow the representation of various shapes, including cylinder , ellipsoids, paraboloids, and hyperboloids. As a result, it generalizes 3D Gaussian primitives [21] and includes 2D Gaussian splatting [16] as a special case, allowing higher-order surface adaptation while incorporating kernel-level innovations such as shape kernels [18], texture billboards [43, 34, 39], and weight distributions [14]. In particular, there is an efficient ray-quadratic intersection algorithm [35] that makes this representation practical and versatile in the context of splatting. Since the primitive is essentially a second-order primitive, we term our method Quadratic Gaussian Splatting (QGS).
However, for deformable quadratic primitives, defining a continuous density distribution is a key challenge. Unlike methods using rigid primitives (e.g., triangles, spheres) with Euclidean-based density, our deformable quadratic surfaces require a geometry-aware density metric. For instance, as Fig 4 shows, when a primitive deforms from a disk to a paraboloid, Euclidean distance would measure straight-line distance, which ignores the curved surface. While Geodesic distance follows the surface, yields more accurate density modeling, it is generally difficult to compute.
@@
Figure 2: Demo and mesh reconstruction comparison on the Mip-NeRF 360 dataset. ellipsoids struggle to align accurately with surfaces, producing coarse and incomplete meshes. Planes fail to capture high-curvature regions, resulting in oversmoothed reconstructions. In contrast, quadric surfaces effectively capture geometric edge regions, achieving detailed reconstruction results.
Luckily, we target at surface modeling so we constrain our quadratic representation to paraboloids—including planar ellipses as degenerate cases—where efficient closed-form geodesic solvers exist. This formulation maintains point-based rendering efficiency while enabling adaptive curvature modeling, where a single paraboloid approximates complex geometry that otherwise requires multiple planar surfels, achieving higher fidelity with similar primitive counts. With such surfel-based representation and inspired by 2DGS [16], we adopt a ray-splat intersection [35] for splatting, maintaining multi-view consistency and perspective correctness. The surfel nature also allow QGS to leverage advanced splatting techniques that address perspective distortion [16] and popping effects [33], or advanced surface regularization [16, 2, 11] that improves overall geometry accuracy, albeit with added computational cost.
To validate the effectiveness, we evaluate QGS on three benchmark datasets—MipNeRF360 [2], DTU [19], and Tanks and Temples [22]—and demonstrate state-of-the-art surface reconstruction accuracy. Our ablation studies reveal that deformable primitives improve geometric fidelity by 33% over 2DGS [16], establishing QGS as a robust alternative for applications demanding higher geometric precision. We summarize our key contributions as follows:
  • A quadric surfel representation that generalizes 2D Gaussians to higher-order surfaces (e.g., paraboloids), enabling adaptive curvature modeling and capturing high frequency details.
  • Geodesic-aware density distributions, the first method to unify deformable primitives with surface-aligned density modeling via closed-form geodesic solvers.
  • State-of-the-art reconstruction: QGS reduces geometric error by 33% over 2DGS and 27% over GOF on DTU [19] while maintaining competitive training and rendering speeds.
2 Related work
2.1 Novel View Synthesis
Novel view synthesis has advanced rapidly, particularly since NeRF [31]. NeRF represents scene geometry and view-dependent appearance using a multi-layer perceptron (MLP) and optimizes it end-to-end through volumetric rendering to generate realistic images. Subsequent methods have introduced various enhancements. Mip-NeRF [1], Mip-NeRF 360 [2], and Zip-NeRF [3] addressed aliasing issues by refining sampling strategies, while I-NGP [32] and DVGO [38] significantly improved training and rendering speeds by replacing MLPs with feature grids.
More recently, 3DGS [21] has gained attention for its superior rendering quality and real-time performance, often surpassing NeRF [49, 33, 30, 54]. Mip-Splatting [49] introduced a smoothing filter to reduce high-frequency artifacts, while StopThePop [33] improved multi-view consistency through per-tile and per-pixel sorting. Scaffold-GS [30] optimized Gaussian distributions by reducing redundancy and improving rendering quality through anchor-based 3D Gaussian placement.
2.2 Neural Surface Reconstruction
Neural volumetric rendering methods often produce smoother and more complete reconstructions than traditional approaches. NeuS [42] and VolSDF [45] introduced signed distance fields (SDF) to model scene geometry within the NeRF framework [31]. Geo-NeuS [11] and NeuralWarp [12] further improved reconstruction by enforcing multi-view consistency. Neuralangelo [24] reduced reliance on MLPs by incorporating feature grids and numerical gradients, enhancing geometric fidelity.
However, neural rendering-based surface reconstruction often takes hours to converge using only image supervision. Additionally, the structured implicit representations make direct access and editing of scene geometry challenging.
2.3 Gaussian Splatting Surface Reconstruction
With the rapid advancement of 3DGS across various domains, Gaussian splatting methods have also significantly improved surface reconstruction. SuGaR [13] introduced regularization terms to encourage Gaussians to align with scene surfaces, constructing volumetric density fields from Gaussian ellipsoids and applying Poisson reconstruction [20] to extract meshes. GSDF [48] and NeuSG [6] combined 3D Gaussian ellipsoids with signed distance fields (SDFs), achieving high-quality geometric reconstructions and rendering results. Other approaches have leveraged ray-Gaussian intersections for improved accuracy. Rade-GS [51], GOF [50], and PGSR [5] computed ray-ellipsoid intersections to obtain unbiased depth estimates and applied normal consistency supervision from 2DGS [16] to achieve state-of-the-art reconstruction. 2DGS further refined this approach by flattening Gaussian ellipsoids into disks, better aligning primitives with surfaces while introducing additional regularization losses for enhanced geometric constraints. MVG-splatting [25] extended 2DGS by incorporating multi-view consistency constraints, enabling more complex surface reconstructions. In this work, we propose Quadratic Gaussian Splatting (QGS), a novel surface representation based on quadric surfaces designed to improve the geometric fitting of primitives. We establish Gaussian distributions on quadrics, enabling end-to-end optimization. Additionally, by integrating multi-view consistency supervision, QGS achieves state-of-the-art geometric reconstruction and high-quality rendering.
3 Method
3.1 Preliminary
Kerbl et al. [21] proposed representing a scene using 3D Gaussian ellipsoids as primitives and render images using differentiable volume splatting.
Here, $\alpha_{i}$ represents the opacity, $c_{i}$ denotes the color of each Gaussian primitive modeled by spherical harmonics, which is also used in our method, and $\mathbf{p}$ is the pixel coordinate. Finally, each Gaussian primitive is optimized by minimizing the photometric loss.
GS mesh reconstruction. Vanilla 3DGS [21] can render high-quality images, but it yields suboptimal results for scene geometry reconstruction. In subsequent surface reconstruction methods, volumetric approaches like GOF [50] and RadeGS [51] employ ray-splat intersection techniques, achieving SOTA reconstruction quality but limiting the consistency of normal and depth. In contrast, 2DGS [16] defines the 2D Gaussian distribution within a planar disk, which inherently provides consistent normals and depths across multiple views. But the disk is only a first-order approximation of the surface, which often leads to overly smooth reconstruction results in 2DGS, as shown in Fig 2.
@@ 3.2 Quadratic Gaussian Splatting
Quadratic Gaussian Model. Given a homogeneous coordinate $\mathbf{x}=[x,y,z,1]^{T}\in\mathbb{R}^{4}$, a quadric surface can be defined as the solution set to the following equation:
@@
We apply a congruent transformation to obtain the transformation from object space to parameter space.
$\mathbf{Q}=\mathbf{T}^{-T}\mathbf{D}\mathbf{T}^{-1},\mathrm{with}$ $\mathbf{D}\;\mathrm{symmetrical}$ (3)
$\mathbf{T}=\left[\begin{matrix}\mathbf{R}&\mathbf{c}\\ 0&1\end{matrix}\right]$
Here, $\mathbf{c}$ denotes the position of the quadric, and $\mathbf{R}=[\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}]$ denotes the orientation of the quadric in the object space. The matrix $\mathbf{D}$ defines the surface scale and shape: $\mathbf{D}=\mathrm{diag}(s_{1}^{2},s_{2}^{2},s_{3}^{2},1)$ yields an ellipsoid, while $\mathbf{D}=\mathrm{diag}(s_{1}^{2},s_{2}^{2},0,0)$ produces a plane.
To compute the Gaussian weight at any surface point, we first define a measure on the surface to establish the Gaussian distribution. A straightforward approach is to construct the Gaussian distribution using Euclidean distance. However, this leads to inconsistencies when the surface undergoes deformation, resulting in uneven splatting reconstruction artifacts, as shown in Fig. 5. To address this, we use geodesic distance to ensure that weights adapt consistently to surface deformations. Fig. 4 shows a 1D illustration.
Figure 4: Density distribution among deformable primitives (disk→paraboloid). (a) Euclidean distance fails to adapt to curvature during primitive deformation, distorting density distributions (uneven weight in the blue point). (b) Our geodesic distance follows intrinsic curvature, maintaining spatial coherence across deformation and preventing uneven splatting artifacts.
However, not all geodesics have closed-form solutions. When $\mathbf{Q}$ is an ellipsoid or hyperboloid, computing the geodesic length typically requires numerical methods. Luckily, as we aim at surface reconstruction, we focus on paraboloid with the following form:
$f(x,y,z)$ $=\mathbf{x}^{T}\left[\begin{matrix}\mathbf{R}&\mathbf{c}\\ 0&1\end{matrix}\right]^{-T}\mathbf{D}\left[\begin{matrix}\mathbf{R}&\mathbf{c}\\ 0&1\end{matrix}\right]^{-1}\mathbf{x}$ (4)
@@
(a) w/ Euclid
(b) w/ Geodesic
Figure 5: Comparison between Euclidean and geodesic distances. The first row presents schematic illustrations, while the second row shows normal comparisons. The left column uses Euclidean distance, while the right column employs geodesic distance.
Gaussian Distribution on Quadric. We will now describe how to define a Gaussian distribution on a paraboloid using geodesic lines. First, paraboloid (Equation 5) could be expressed in explicit form:
@@
$\begin{split}\hat{z}(\theta,\rho)=\;&s_{3}(\frac{\mathrm{sign}(s_{1})\cos^{2}\theta}{s_{1}^{2}}+\frac{\mathrm{sign}(s_{2})\sin^{2}\theta}{s_{2}^{2}})\rho^{2}\\ =\;&a(\theta)\rho^{2}\end{split}$ (7)
@@
(a) normal
(b) curvature
Figure 6: The (a) normal map and (b) curvature map of a scene. In the curvature map, blue indicates higher curvature and curved areas, while red indicates lower curvature and flatter areas.